Sandra's brother is 6 years older than her, her sister is 2 years younger than her, and her mother is 3 times her age. Her brother, sister, and mother's combined age is 64. Write an equation to solve for Sandra's age.
Answer by sachi(548)

let sandra is x yrs old
than her brother is x+6
her sister is x-2
& her mother is 3x
now her brother, sister, and mother's combined age is 64
so x+6+x-2+3x=64
or 6x=64-4=60
or x=10
so sandra is 10 yrs old,her brother is x+616, her sister is 8 & her mother is 30 yrs
ans